# Finite Measures of Maximal Entropy for an Open Set of Partially Hyperbolic Diffeomorphisms

### v1

@@ Abstract.
We consider partially hyperbolic diffeomorphisms $f$ with a one-dimensional central direction such that the unstable entropy exceeds the stable entropy. Our main result proves that such maps have a finite number of ergodic measures of maximal entropy. Moreover, any $C^{1+}$ diffeomorphism near $f$ in the $C^{1}$ topology possesses at most the same number of ergodic measures of maximal entropy. Our contribution is in extending the findings in [2] to arbitrary dimensions. We believe our technique, essentially distinct from the one in that paper, is robust and may find applications in further contexts.
@@ 1.1. Measures of maximal entropy
A compelling feature of dynamical systems involves the rate at which the number of distinguishable orbits grows, a quantifiable aspect captured by the concept of entropy. The computation of entropy can be examined through both topological and invariant measure lenses, interconnected by the variational principle. This principle offers a method to determine the topological entropy of a dynamical system by exploring its invariant measures and the entropies associated with them.
To be more precise, let $f:M\to M$ be a continuous function on a compact metric set. Denoting by ${\mathcal{M}}_{f}$ the set of $f$-invariant probability measures, the variational principle establishes the following relation
@@
Sinai, Ruelle and Bowen pioneered the study of measures of maximal entropy in the 1970s. Sinai, a trailblazer in the field, established the existence and finiteness of maximal entropy measures for Anosov diffeomorphisms [31]. Subsequently, Ruelle and Bowen extended and broadened this approach to encompass uniformly hyperbolic (Axiom A) systems [3, 27, 28].
Newhouse demonstrated that $C^{\infty}$ diffeomorphisms on a compact manifold exhibit maximum entropy measures [20]. He posed the question of whether the count of such measures is finite, particularly for diffeomorphisms on surfaces with positive topological entropy [21, Problem 2].
As the skew product of a linear Anosov diffeomorphism by a rational rotation exhibits an infinite number of such measures, this question is relevant only for surfaces. Recently, Buzzi, Crovisier and Sarig provided a positive answer to this question in [2].
In our previous work [18], we delved into the characteristics of unstable and stable entropy, denoted as $h^{u}(f)$ and $h^{s}(f)$ respectively, as introduced in [33, 10]. By employing the criterion developed in [7, 25], we established that partially hyperbolic systems possessing a unidimensional central bundle, wherein the unstable entropy surpasses the stable entropy and the unstable foliation is minimal, there robustly exists a unique measure of maximal entropy.
In the present work, we withdraw the hypotheses about the minimality of the unstable foliation and prove:
Main Theorem. Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$, then $f$ has a finite number of ergodic measures of maximal entropy. Moreover, there exists a $C^{1}$ neighborhood $U$ of $f$, such that for any $C^{1+}$ diffeomorphism $g$ within $U$, the number of ergodic measures of maximal entropy of $g$ is not greater than the number of maximal entropy ergodic measures of $f$.
@@ 1.2. Homoclinic classes with high entropy
The homoclinic class of a hyperbolic periodic orbit of saddle type ${{\mathcal{O}}}$ is defined by
@@
An ergodic invariant probability measure $\mu$ is hyperbolic of saddle type, if it has at most one positive Lyapunov exponent, at most one negative Lyapunov exponent, and no zero Lyapunov exponent. When $\mu$ is a hyperbolic measure for $f$ and ${{\mathcal{O}}}\in{{\rm Per}}_{h}(f)$ we write $\mu\sim{{\mathcal{O}}}$ when $\mu$ is homoclinically related to ${{\mathcal{O}}}$, see Subsection 2.2, definitions 2.3 and 2.6.
The next theorem establishes that for a diffeomorphism $f$ as described in the Main Theorem, there exists a finite set of hyperbolic periodic orbits such that every ergodic measure with high entropy is related to one of them.
@@
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$, then for any $h(f)\geq a>h^{s}(f)$ there exists an integer $K_{a}>0$ and $K_{a}$ hyperbolic orbits ${{\mathcal{O}}}_{1},\cdots{{\mathcal{O}}}_{K_{a}}$ with $s-$index ${\operatorname{dim}}(E^{s})+1$ such that $h(f_{|HC({{\mathcal{O}}}_{i})})\geq a$ for every $i=1,\cdots K_{a}$ and any $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(g)>a$, $\mu\sim{{\mathcal{O}}}_{i}$ for some $1\leq j\leq K_{a}$.
Remark 1.2.
It is important to note that the above theorem does not imply a spectral decomposition theorem like [2, Theorem 1]. Our main difficulty arises from the potential existence of two hyperbolic periodic orbits that are not homoclinically related, but one is entirely contained within the other. Consequently, it is impossible to make any statement about the entropy of the intersection between two homoclinic classes in Theorem A.
Remark 1.3.
The approach used by Buzzi, Crovisier and Sarig in their notable work in [2] establishes that $C^{\infty}$ surface diffeomorphisms with positive topological entropy have at most a finite number of ergodic measures of maximum entropy in general, and at most one in the case topologically transitive, has four steps. The first step heavily uses two-dimensionality to bind topological disks ($su$-quadrilaterals) by two stable and two unstable segments. This allows them to obtain topological intersections between stable and unstable manifolds of two different measures. In the second step, taking advantage of the fact that the diffeomorphism under consideration has high entropy and a dynamic version of a lemma by Sard, they conclude that some of these intersections are transversal. The third step consists of proving the uniqueness of m.m.e. between a set of hyperbolic measures that are homoclinically related. In the fourth step, using the properties of su-quadrilaterals in step one, they prove that two hyperbolic measures carried by a common transitive set have intersections from their stable and unstable foliations. Applying the results of the previous steps, they were able to prove that any stable leaf of one measure must transversally intersect some leaf of the unstable foliation of the other measure, leading to a homoclinic relation between the measures. Thus, the technique used in [2] is strongly 2-dimensional and we can not repeat it in a higher dimension.
Remark 1.4.
In general, it does not hold that every diffeomorphism close to a diffeomorphism with a finite number of measures of maximal entropy has also a finite number of measures of maximal entropy. But for diffeomorphisms as in the theorem above we prove that every $C^{1+}$ diffeomorphism sufficiently close to $f$ also has a finite number of ergodic m.m.e., see Corollary 5.2.
@@ 1.3. Dependence of the number of ergodic measures of maximal entropy on the diffeomorphism
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$. The function $N_{max}:\operatorname{Diff}^{1+}(M)\to\mathbb{N}\cup\{\infty\}$ is upper semicontinuous in the $C^{1}$ topology, that is, $\limsup_{g\to f}N_{max}(g)\leq N_{max}(f)$.
@@ 1.4. Equilibrium states
Theorem C.
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$, ${\operatorname{dim}}(E^{c})=1$ and $\phi$ a potential Hölder continuous. If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, then for any $P(f,\phi)\geq a>h^{s}(f)+\sup\phi$ there exists an integer $K_{a}>0$ and $K_{a}$ hyperbolic orbits $\mathcal{O}_{1},\cdots\mathcal{O}_{K_{a}}$ with $s-$index ${\operatorname{dim}}(E^{s})+1$ such that $P(f_{|HC({{\mathcal{O}}}_{i})},\phi_{|HC({{\mathcal{O}}}_{i})})\geq a$ for every $i=1,\cdots K_{a}$ and any $\mu\in{\mathcal{M}}_{f}^{e}$ with $P_{\mu}(f,\phi)\geq a$, $\mu\sim\mathcal{O}_{i}$ for some $1\leq j\leq K_{a}$.
As a consequence of Theorem C, we get that $C^{1+}$ diffeomorphisms in a $C^{1}$ open neighborhood of $f$ also have a finite number of equilibrium states, see Corollary 6.2.
Let $M_{\text{max}}(f)$ be the number of ergodic equilibrium states of a diffeomorphism $f$. Our last result establishes the following upper semicontinuity property for $M_{\text{max}}(f)$.
@@
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$. If $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, The function $M_{max}:{{\rm Diff}}^{1+}(M)\to{\mathbb{N}}\cup\{\infty\}$ varies upper semicontinuity in the $C^{1}$ topology, that is, $\limsup_{g\to f}M_{max}(g)\leq M_{max}(f)$.
1.5. Structure of the paper and the sketch of the proof of the Main Theorem

### v3

@@ Abstract.
We consider partially hyperbolic diffeomorphisms $f$ with a one-dimensional central direction such that the unstable entropy is different from the stable entropy. Our main result proves that such maps have a finite number of ergodic measures of maximal entropy. Moreover, any $C^{1+}$ diffeomorphism near $f$ in the $C^{1}$ topology possesses at most the same number of ergodic measures of maximal entropy. These results extend the findings in [4] to arbitrary dimensions and provides an open class of non Axiom A systems of diffeomorphisms exhibiting a finite number of ergodic measures of maximal entropy. We believe our technique, essentially distinct from the one in [4], is robust and may find applications in further contexts.
@@ 1. Introduction
A compelling feature of dynamical systems involves the rate at which the number of distinguishable orbits grows, a quantifiable aspect captured by the concept of entropy. The computation of entropy can be examined through both topological and invariant measure lenses, interconnected by the variational principle: the topological entropy is the supremum over the metric entropies of all invariant probability measures. An invariant measure that attains the highest entropy is called a measure of maximal entropy and its study is a natural way of describing the behavior of most of the relevant orbits of a system.
Sinai, Margulis, Ruelle and Bowen pioneered the study of measures of maximal entropy in the 1970s. Sinai and Margulis, the trailblazers in the field, established the existence and finiteness of maximal entropy measures for Anosov diffeomorphisms [48, 26]. Subsequently, Ruelle and Bowen extended and broadened this approach to encompass uniformly hyperbolic (Axiom A) systems [6, 43, 44].
In [32] Newhouse demonstrated that $C^{\infty}$ diffeomorphisms on a compact manifold exhibit maximal entropy measures (but finite smoothness does not, according to Misiurewicz [29]). Newhouse posed the question of whether the number of such measures is finite, particularly for diffeomorphisms on surfaces with positive topological entropy [33, Problem 2]. As the skew product of a linear Anosov diffeomorphism by a rational rotation exhibits an infinite number of such measures, this question is relevant only for surfaces. Recently, Buzzi, Crovisier and Sarig provided a positive answer to this question in [4].
In the context of partially hyperbolic diffeomorphisms, the existence and uniqueness of equilibrium states is not clear at all. In three dimensional manifolds, Buzzi et all proved in [5] that a certain derived from Anosov construction due to Mañé also has this uniqueness property. In [34], Newhouse and Young have shown that some partially hyperbolic examples on $T^{4}$ have a unique measure of maximal entropy. The results in [13] includes the uniqueness of ergodic measure of maximal entropy for Mañé’s example in [27]. Recently, in [47], the authors prove that Shub’s example has a unique measure of maximal entropy.
For accessible partially hyperbolic diffeomorphisms of 3-dimensional manifolds having compact center leaves, either there is a unique measure of maximal entropy, or there are a finite number of hyperbolic measures of maximal entropy, see [41]. In [50] the authors characterize the maximal entropy measure for $C^{2}$ partially hyperbolic systems with circle central bundles by means of suitable finite set of saddle points.
Let us also mention that establishing the existence and, more importantly, the uniqueness of ergodic maximal entropy measures has been successful for flows without any singularity. Among such works, let us mention the works [8, 21, 3, 1, 16, 47].
To establish the number of maximum entropy measures for flows with singularities is a challenging task. To the best of the authors’ knowledge, there is only one article currently available on this subject, [38], where it is proved that for a sectional-hyperbolic attractor $\Lambda$, in a $C^{1}$-open and dense family of flows (including the classical Lorenz attractor, [25]), if the point masses at singularities are not equilibrium states, then there exists a unique equilibrium state supported on $\Lambda$. In particular, there exists a unique measure of maximal entropy for the flow restricted to $\Lambda$.
Back to the diffeomorphism context, we recall that the notion of unstable and stable entropy, denoted as $h^{u}(f)$ and $h^{s}(f)$ respectively, was introduced in [19], reminiscent of the works by Yomdin and Newhouse ([54, 32]) and studied in [53, 17, 49].
In [51], it is shown that there is a finite number of maximal u-entropy measures for $C^{1}$ diffeomorphisms that factor over Anosov and has mostly contracting center. In the same setting as in [51], the same authors prove the exponential decay of correlations for a u-measure of maximal entropy [52].
In our previous work [30], we delved into the characteristics of unstable and stable entropy. By employing the criterion developed in [11, 37], we established that for partially hyperbolic systems possessing a unidimensional central bundle, wherein the unstable entropy surpasses the stable entropy and the unstable foliation is minimal, there robustly exists a unique ergodic measure of maximal entropy.
In this paper, we address the problem of the existence and uniqueness (finiteness) of entropy maximal measures in the partially hyperbolic context. We consider partially hyperbolic diffeomorphisms $f$ with a one-dimensional central direction such that the unstable entropy exceeds the stable entropy. Our main result proves that such maps have a finite number of ergodic measures of maximal entropy. Moreover, any $C^{1+}$ diffeomorphism near $f$ in the $C^{1}$ topology possesses at most the same number of ergodic measures of maximal entropy. These results extend the findings in [4] to higher dimensions and provides an open class of non Axiom A systems of diffeomorphisms exhibiting a finite number of ergodic measures of maximal entropy. To be more precise, let $f:M\to M$ be a continuous function on a compact metric set. Denoting by ${\mathcal{M}}_{f}$ the set of $f$-invariant probability measures, the variational principle establishes the following relation
@@
In the present work, we withdraw the hypotheses about the minimality of the unstable foliation in [30] and prove:
Main Theorem. Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$. If $h^{u}(f)\neq h^{s}(f)$, then $f$ has only a finite number of ergodic measures of maximal entropy. Moreover, there exists a $C^{1}$ neighborhood $U$ of $f$, such that for any $C^{1+}$ diffeomorphism $g$ within $U$, the number of ergodic measures of maximal entropy of $g$ is not greater than the number of maximal entropy ergodic measures of $f$.
@@ 1.1. Homoclinic classes with high entropy
The homoclinic class of a hyperbolic periodic orbit ${{\mathcal{O}}}$ of saddle type is defined by
@@
An ergodic invariant probability measure $\mu$ is hyperbolic of saddle type, if it has at least one positive Lyapunov exponent, at least one negative Lyapunov exponent, and no zero Lyapunov exponent. When $\mu$ is a hyperbolic measure for $f$ and ${{\mathcal{O}}}\in{{\rm Per}}_{h}(f)$ we write $\mu\sim{{\mathcal{O}}}$ when $\mu$ is homoclinically related to ${{\mathcal{O}}}$, see Subsection 2.2, definitions 2.3 and 2.6.
The next theorem establishes that for a diffeomorphism $f$ as described in the Main Theorem, there exists a finite set of hyperbolic periodic orbits such that every ergodic measure with high entropy is homoclinically related to one of them.
Since $h^{u}(f)=h^{s}(f^{\,-1}),$ see Definition 2.14, from now on, we only consider the case $h^{u}(f)>h^{s}(f)$; the other case follows by considering $f^{\,-1}$.
@@
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a close manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$, then for any $h(f)\geq a>h^{s}(f)$ there exists an integer $K_{a}>0$ and $K_{a}$ hyperbolic orbits ${{\mathcal{O}}}_{1},\cdots{{\mathcal{O}}}_{K_{a}}$ with $\operatorname{s-index}\,{\operatorname{dim}}(E^{s})+1$ such that $h(f_{|HC({{\mathcal{O}}}_{i})})\geq a$ for every $i=1,\cdots K_{a}$ and any $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(g)>a$ must satisfy $\mu\sim{{\mathcal{O}}}_{i}$ for some $1\leq j\leq K_{a}$.
It is important to note that the above theorem does not imply a spectral decomposition theorem like [4, Theorem 1]. Our main difficulty arises from the potential existence of two hyperbolic periodic orbits that are not homoclinically related, but their homoclinic classes are such that one is contained in the other. Consequently, it is impossible to make any statement about the entropy of the intersection between two homoclinic classes in Theorem A.
Buzzi, Crovisier and Sarig in their notable work in [4] establishes that $C^{\infty}$ surface diffeomorphisms with positive topological entropy have at most a finite number of ergodic measures of maximum entropy in general. The approach there is strongly supported by the two-dimensionality to bind topological disks ($su$-quadrilaterals) by two stable and two unstable segments. These $su$-quadrilaterals allow them to obtain topological intersections between stable and unstable manifolds of two different measures with large entropy that are weak close. Using a dynamic version of a lemma by Sard, they conclude that some of these intersections are transversal leading to a homoclinic relation between the measures. Therefore, replicating this strategy in a higher-dimensional context is unfeasible. We, instead, propose an alternative method in which we define a set where each of its points has a stable Pesin manifold of uniform size. Taking advantage of Pliss’s Lemma, we then prove that the measure of this set for measures with high entropy is strictly positive. Then we were able to prove that for nearby points in this set their stable and unstable manifolds intersect transversaly. This implies that the number of homoclinic classes with measures with large entropy is finite. And then, using [4, Corollary 2.14], we conclude that the number of ergodic maximal entropy measures of $f$ is finite. We believe that our method can be applied to other higher dimensional contexts.
In general, it does not hold that every diffeomorphism close to a diffeomorphism with a finite number of measures of maximal entropy has also a finite number of measures of maximal entropy. But for diffeomorphisms as in the theorem above we prove that every $C^{1+}$ diffeomorphism sufficiently close to $f$ also has only a finite number of ergodic m.m.e., see Corollary 5.2.
@@ 1.2. Dependence of the number of ergodic measures of maximal entropy on the diffeomorphism
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. If $h^{u}(f)>h^{s}(f)$ then the function $N_{max}:\operatorname{Diff}^{1+}(M)\to\mathbb{N}\cup\{\infty\}$ is upper semicontinuous in the $C^{1}$ topology, that is, $\limsup_{g\to f}N_{max}(g)\leq N_{max}(f)$.
The meaning of this theorem lies in resolving a long-standing question within smooth dynamical systems. Although Bowen established the robustness of the number of ergodic measures of maximum entropy for Axiom A diffeomorphisms, the search for non-Axiom A examples remained an open challenge. This theorem provides a non-Axiom A system that maintains the property of finiteness of ergodic measures of maximal entropy under small perturbations. In addition, the quantity of such measures of these perturbations are bounded by the number of ergodic maximal entropy measures of $f$. The proof of this theorem involves a refined analysis of the properties of unstable and stable entropy, as well as certain estimates related to a set of points exhibiting a stable manifold of uniform size.
1.3. Central Lyapunov exponent far from zero
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. Consider the sets
${\mathcal{M}}^{+}_{f}:=\{\mu\in{\mathcal{M}}_{f}^{e}:\lambda_{c}(f,\mu)\geq 0\}\quad\mbox{ and }\quad{\mathcal{M}}^{-}_{f}:=\{\mu\in{\mathcal{M}}_{f}^{e}:\lambda_{c}(f,\mu)\leq 0\}$
and let
$h^{-}(f):=\sup_{\mu\in{\mathcal{M}}_{f}^{-}}\{h_{\mu}(f)\}\quad\mbox{ and }\quad h^{+}(f):=\sup_{\mu\in{\mathcal{M}}_{f}^{+}}\{h_{\mu}(f)\}.$
In [12], it was proved that a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with one-dimensional center, dense stable and unstable foliations, and $h^{+}(f)\neq h^{-}(f)$ has a unique measure of maximal entropy. We prove the following theorem.
Theorem C.
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and $\operatorname{dim}(E^{c})=1$. If $h^{+}(f)\neq h^{-}(f)$ then $f$ has only a finite number of ergodic maximal entropy.
The condition $h^{-}(f)>h^{+}(f)$ is equivalent to requiring that every ergodic measure with large entropy has a central Lyapunov exponent far from zero, as stated in Proposition 6.1.
When $f$ satisfies the conditions of Theorem A, we can conclude that $h^{-}(f)>h^{+}(f)$. However, it is not clear whether $h^{-}(f)>h^{+}(f)$ implies $h^{u}(f)>h^{s}(f)$. Therefore, we cannot obtain information about diffeomorphisms close to $f$ when $f$ satisfies $h^{-}(f)>h^{+}(f)$.
@@ 1.4. Equilibrium states
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$, ${\operatorname{dim}}(E^{c})=1$ and $\phi$ a Hölder continuous potential. If green $h^{u}(f)>\sup\phi\geq\inf\phi>h^{s}(f)$, then for any $P(f,\phi)\geq a>h^{s}(f)+\sup\phi$ there exists an integer $K_{a}>0$ and $K_{a}$ hyperbolic orbits $\mathcal{O}_{1},\cdots\mathcal{O}_{K_{a}}$ with $s-$index ${\operatorname{dim}}(E^{s})+1$ such that $P(f_{|HC({{\mathcal{O}}}_{i})},\phi_{|HC({{\mathcal{O}}}_{i})})\geq a$ for every $i=1,\cdots K_{a}$ and any $\mu\in{\mathcal{M}}_{f}^{e}$ with $P_{\mu}(f,\phi)\geq a$ must satisfy $\mu\sim\mathcal{O}_{i}$ for some $1\leq j\leq K_{a}$.
As a consequence of Theorem D, we get that $C^{1+}$ diffeomorphisms in a $C^{1}$ open neighborhood of $f$ also have a finite number of equilibrium states, see Corollary 7.2.
Let $M_{\text{max}}(f,\varphi)$ be the number of ergodic equilibrium states of a diffeomorphism $f$ and a potential $\varphi$. Our last result establishes the following upper semicontinuity property for $M_{\text{max}}(\cdot,\varphi)$.
Theorem E.
Let $f:M\to M$ be a $C^{1+}$ partially hyperbolic diffeomorphism of a compact manifold $M$ with $TM=E^{u}\oplus E^{c}\oplus E^{s}$ and ${\operatorname{dim}}(E^{c})=1$, and $\phi$ a Hölder continuous potential. If $h^{u}(f)>\sup\phi\geq\inf\phi>h^{s}(f)$, then the function ${}M_{max}(.,\phi):{{\rm Diff}}^{1+}(M)\to{\mathbb{N}}\cup\{\infty\}$ varies upper semi-continuously in the $C^{1}$ topology, that is, $\limsup_{g\to f}M_{max}(g)\leq M_{max}(f)$.
1.5. Sketch of the proof of the Main Theorem